# Fast Frank–Wolfe Algorithms with Adaptive Bregman Step-Size for Weakly Convex Functions

### v3

@@ Abstract
We propose a Frank–Wolfe (FW) algorithm with an adaptive Bregman step-size strategy for smooth adaptable (also called: relatively smooth) (weakly-) convex functions. This means that the gradient of the objective function is not necessarily Lipschitz continuous, and we only require the smooth adaptable property. Compared to existing FW algorithms, our assumptions are less restrictive. We establish convergence guarantees in various settings, such as sublinear to linear convergence rates, depending on the assumptions for convex and nonconvex objective functions. Assuming that the objective function is weakly convex and satisfies the local quadratic growth condition, we provide both local sublinear and local linear convergence regarding the primal gap. We also propose a variant of the away-step FW algorithm using Bregman distances over polytopes. We establish global faster (up to linear) convergence for convex optimization under the Hölder error bound condition and its local linear convergence for nonconvex optimization under the local quadratic growth condition. Numerical experiments demonstrate that our proposed FW algorithms outperform existing methods.
@@ 1 Introduction
where $f:\mathbb{R}^{n}\to(-\infty,+\infty]$ is a continuously differentiable function and $P\subset\mathbb{R}^{n}$ is a compact convex set. We are interested in both convex and nonconvex $f$ and assume that we have first-order oracle access to $f$, i.e., given $x\in\mathbb{R}^{n}$, we can compute $\nabla f(x)$. The Frank–Wolfe (FW) algorithm is a projection-free first-order method. Instead of requiring access to a projection oracle, the FW algorithm requires only access to a so-called linear minimization oracle (LMO), which for a given linear function $a\in\mathbb{R}^{n}$ computes $y\in\operatorname*{argmin}_{v\in P}\langle a,v\rangle$. LMOs are often much cheaper than projection oracles, as shown in [CP2021] (see also [Braun2022-ew, Table 1.1]), so that, in practice, FW algorithms are often faster than projected gradient methods even when the projection operation is not trivial. Additionally, FW algorithms tend to be numerically quite robust and stable due to their affine-invariance and can be also used, e.g., to provide theoretical guarantees for the approximate Carathéodory problem [cp2019approxCara].
@@
The FW algorithm was originally proposed by Frank and Wolfe [Frank1956-cq] and was independently rediscovered and extended by Levitin and Polyak [Levitin1966-sm] as the conditional gradient method; we will use these terms interchangeably. Canon and Cullum [Canon1968-sr] established an initial lower bound for the convergence rate of the FW algorithm, which was later improved by GuéLat and Marcotte [GueLat1986-bg], which also provided the first analysis of the away-step FW algorithm by Wolfe [Wolfe1970-ax]. Jaggi [Jaggi2013-oz] provided a more detailed convergence analysis of the FW algorithm, establishing a new lower bound that demonstrated a trade-off between sparsity and error. Concurrently, Lan [lan2013complexity] examined the complexity of linear programming-based first-order methods, establishing a similar lower bound. For a comprehensive review of FW algorithms, we refer the interested reader to the survey by Braun et al. [Braun2022-ew] and the brief introduction by Pokutta [Pokutta2024-qu].
The classical FW algorithm (Algorithm 1) has been improved in two main directions to enhance both performance and convergence. One focuses on refining the step-size rule. The short-step strategy computes $\gamma_{t}$ using the Lipschitz constant $L$ of $\nabla f$. Pedregosa et al. [Pedregosa2020-ay] introduced an adaptive step-size strategy that dynamically estimates $L$. They proved that this strategy is at least as effective as the short-step strategy. It was enhanced by Pokutta [Pokutta2024-qu] to improve numerical stability.
Another direction is to modify the classical FW algorithm to eliminate the zigzag behavior when approaching the optimal face containing the optimal solution $x^{*}$. This behavior inspired Wolfe [Wolfe1970-ax] to propose the away-step FW algorithm in 1970 that shortcuts the zigzagging by removing atoms that slow down the iterative sequence. Lacoste-Julien and Jaggi [Lacoste-Julien2015-gb] showed the linear convergence of the away-step FW algorithm. Their analysis***Similar analyses based on the pyramidal width have been developed for other advanced variants, such as the pairwise FW algorithm [Lacoste-Julien2015-gb], Wolfe’s algorithm [Lacoste-Julien2015-gb], the blended conditional gradient method [BPTW2018], and blended pairwise conditional gradient methods and similar variants [CP2020boost, TTP2021]. Even Nesterov-style acceleration is possible under weak assumptions [CDP2019] building upon this analysis. introduced a geometric constant, the pyramidal width, that measures the conditioning of the polytope $P$, representing the feasible region.
@@
To derive the convergence rate of the above-mentioned FW algorithms, previous studies have typically required $L$-smoothness, i.e., that $\nabla f$ is Lipschitz continuous, and the convexity for the objective function $f$, although exceptions exist.Quite recently, Vyguzov and Stonyakin [Vyguzov2025-ha] proposed a FW algorithm with the Bregman distance, similar to ours, yet it is restricted to convex problems and provides only partial convergence guarantees. These assumptions often narrow the application of FW algorithms. Even if they do not satisfy $L$-smoothness, there are many functions that satisfy the $L$-smooth adaptable property ($L$-smad) with kernel generating distances $\phi$ (see Definition 2.1) by choosing $\phi$ well to match the function $f$. Since the $L$-smad property is consistent with $L$-smoothness when $\phi=\frac{1}{2}\|\cdot\|^{2}$, the class of $L$-smad functions includes the $L$-smooth function class. Other first-order algorithms, such as the Bregman proximal gradient algorithm, are often analyzed under the $L$-smad property (see e.g., [Bolte2018-zt, Hanzely2021-sz, Rebegoldi2018-pa, Takahashi2022-ml, Takahashi2024-wz, Yang2025-ls]). $L$-smad functions appear in many applications, such as nonnegative linear inverse problems [Bauschke1997-vk, Takahashi2024-wz], $\ell_{p}$ loss problems [Kyng2015-uo, Maddison2021-nt], phase retrieval [Bolte2018-zt, Takahashi2022-ml], nonnegative matrix factorization (NMF) [Mukkamala2019-mk, Takahashi2024-cu], and blind deconvolution [Takahashi2023-uh]. Any $\mathcal{C}^{2}$ function is locally $L$-smooth over compact sets, but the resulting Lipschitz constant $L$ can be overly conservative, hindering practical performance. Functions outside $\mathcal{C}^{2}$ may not satisfy $L$-smoothness at all. Moreover, functions in the above-mentioned applications are weakly convex on compact sets. For nonconvex functions, although Lacoste-Julien [Lacoste-Julien2016-mj] established sublinear convergence, little else has been done, and simply tracing existing research does not achieve convergence rates better than sublinear convergence. Then the following question naturally arises:
Is it possible to loosen the $L$-smoothness and convexity assumptions often made in FW algorithm studies while still achieving the theoretical guarantee of linear convergence rates?
@@
The answer to the above question is yes. We develop FW algorithms for $L$-smad $f$ and give theoretical guarantees of linear convergence under weaker assumptions (Hölder error bound (HEB) condition with parameter $q\geq 1$ for convex $f$ and the special case with $q=2$ called quadratic growth for nonconvex $f$) than the strong convexity assumption that is often made when linear convergence is shown for FW algorithms. The convexity of $f$ is also weaken to weakly convex, i.e., $f+\frac{\rho}{2}\|\cdot\|^{2}$ is convex with some $\rho>0$ (see, e.g.[Davis2018-bb, Examples 3.1 and 3.2] and [Nurminskii1975-vp]). Moreover, it is easy to see that any $\mathcal{C}^{2}$ function defined on a compact set is weakly convex [Vial1983-oo, Proposition 4.11] and [Wu2007-dt]; since $\rho$ is only needed for the theoretical guarantee and not for the algorithm, there is no need for its estimation, and considering weakly-convex $f$ greatly expands the range of applications.
@@
Table 1 summarizes our contributions for the various cases that we consider. Relaxing the assumption from $L$-smoothness to the $L$-smad property makes the construction of a FW algorithm and its theoretical guarantee much more difficult. This is because we use the Bregman distance, which is an extension of the Euclidean distance, for the $L$-smad property. While (2.1) is simple with $\nu=1$ when using the Euclidean distance $\phi=\frac{1}{2}\|\cdot\|^{2}$, $\nu>0$ cannot be eliminated for the general Bregman distance; $1+\nu$ represents its scaling exponent. $\nu$ must be estimated to use in the algorithm since its exact value is unknown. Therefore, the extension to an $L$-smad $f$ is not straightforward.
@@
  • We establish sublinear convergence for the case of convex $L$-smad functions. Assuming that $f$ is convex, satisfies the HEB condition, and the minimizer $x^{*}\in\operatorname{int}P$, we provide the faster (up to linear) convergence. In this setting, our algorithm always converges linearly if the exponent of the HEB condition $q$ equals the scaling exponent of the Bregman distance $1+\nu$, i.e., $q=1+\nu$.As discussed in Sec. 2, linear convergence with the Euclidean distance for convex optimization is shown in [garber2015faster] under the quadratic growth over strongly convex sets. This corresponds to the $q=1+\nu$ case with $\nu=1$. For $q>1+\nu$, if $t\leq t_{0}$ with some $t_{0}\in\mathbb{N}$, linear convergence holds; otherwise, $\mathcal{O}(\epsilon^{(1+\nu-q)/\nu q})$, which is faster than existing sublinear rates. A similar argument also applies to the nonconvex case ($q=2$) and the away-step FW algorithm.
@@
  • For nonconvex optimization, assuming that $f$ is weakly convex and satisfies the local quadratic growth condition, we establish local linear convergence. To our knowledge, this is the first proposal of a FW algorithm with linear convergence guarantees for a certain type of nonconvex optimization problem, albeit of a limited type. These results are also new for the standard setting: $L$-smooth $f$. Without weak convexity, we prove a global sublinear convergence to a stationary point of (1.1) for nonconvex $L$-smad $f$.
@@
Table 1: Convergence rates for Frank–Wolfe algorithms. Our convergence rates are generalizations of existing rates. Moreover, our convergence rate is better than $\mathcal{O}(\epsilon^{-1})$ in some cases. For nonconvex functions, convergence is measured using the Frank–Wolfe gap $\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle\leq\epsilon$, instead of the primal gap $f(x_{t})-f^{*}\leq\epsilon$. Convergence rates for the weakly convex optimization locally hold.
Assumptions Convergence rate
FW $f$ convexity $f$ growth $x^{*}\in\operatorname{int}P$ $P$ polytope $L$-smooth $L$-smad ($\nu>0$)
Alg.1 (any step-size) convex $\mathcal{O}(\epsilon^{-1})$ [Jaggi2013-oz] $\mathcal{O}(\epsilon^{-1/\nu})$ ([Vyguzov2025-ha], Thm.C.1)
Alg.1 (short & adapt.) convex $q$-HEB $\mathcal{O}(\log\epsilon^{-1})$11footnotemark: 1[KDP2018jour] $\mathcal{O}(\log\epsilon^{-1})$ [^2] It holds if the exponent of the HEB condition $q$ equals the scaling exponent of the Bregman distance $1+\nu$, i.e., $q=1+\nu$ ($(\nu,q)=(1,2)$ for the away-step FW algorithm). For $q>1+\nu$, it also holds if $t\leq t_{0}$ with some $t_{0}\in\mathbb{N}$; otherwise $\mathcal{O}(\epsilon^{(1+\nu-q)/\nu q})$.(Thm. 4.2)
away-step FW convex $q$-HEB $\mathcal{O}(\log\epsilon^{-1})$11footnotemark: 1[KDP2018jour] $\mathcal{O}(\log\epsilon^{-1})$22footnotemark: 2(Thm. 4.4)
Alg.1 (any step-size) weak 2-HEB $\mathcal{O}(\epsilon^{-1})$ $\mathcal{O}(\epsilon^{-1/\nu})$ (Thm. D.3)
Alg.1 (short & adapt.) weak 2-HEB $\mathcal{O}(\log\epsilon^{-1})$ $\mathcal{O}(\log\epsilon^{-1})$ [^4] It holds if $\nu=1$. For $\nu\in(0,1)$, it also holds if $t\leq t_{0}$ with some $t_{0}\in\mathbb{N}$; otherwise $\mathcal{O}(\epsilon^{(\nu-1)/2\nu})$.(Thm. 5.2)
away-step FW weak 2-HEB $\mathcal{O}(\log\epsilon^{-1})$ $\mathcal{O}(\log\epsilon^{-1})$44footnotemark: 4(Thm. 5.3)
Alg.1 (any step-size) $\mathcal{O}(\epsilon^{-2})$ [Lacoste-Julien2016-mj] $\mathcal{O}(\epsilon^{-1-1/\nu})$ (Thm. D.1)
@@ 2 Preliminaries
[…]
Let $C$ be a nonempty open convex subset of $\mathbb{R}^{n}$.
Definition 2.1 (Kernel Generating Distance [Bolte2018-zt, Definition 2.1]).
A function $\phi:\mathbb{R}^{n}\to(-\infty,+\infty]$ is called a kernel generating distance associated with $C$ if it satisfies the following conditions:
  1. (i)
    $\phi$ is proper, lower semicontinuous, and convex, with $\operatorname{dom}\phi\subset\operatorname{cl}C$ and $\operatorname{dom}\partial\phi=C$;
  2. (ii)
    $\phi$ is $\mathcal{C}^{1}$ on $\operatorname{int}\operatorname{dom}\phi\equiv C$.
We denote the class of kernel generating distances associated with $C$ by $\mathcal{G}(C)$.
Definition 2.2 (Bregman Distance [Bregman1967-rf]).
Given $\phi\in\mathcal{G}(C)$, a Bregman distance $D_{\phi}:\operatorname{dom}\phi\times\operatorname{int}\operatorname{dom}\phi\to\mathbb{R}_{+}$ associated with $\phi$ is defined by $D_{\phi}(x,y)=\phi(x)-\phi(y)-\langle\nabla\phi(y),x-y\rangle\geq 0$.
We use $D_{f}$ in Definition 2.2 for nonconvex $f$ instead of convex $\phi$ for brevity. In the remainder of the paper, if not stated otherwise, we will assume that as such for a given strictly convex $\phi$ and $C$, there exists $\nu>0$ such that the following holds for all $x,y\in\operatorname{int}\operatorname{dom}\phi$ and $\gamma\in[0,1]$ (see also Lemma A.3, which shows that a weaker one holds)):
@@
We will now recall the smooth adaptable property, which is a generalization of $L$-smoothness and was first introduced by [Bauschke2017-hg]. The smooth adaptable property is also called relative smoothness [Lu2018-ii].
Definition 2.3 ($L$-smooth Adaptable Property [Bolte2018-zt, Definition 2.2]).
@@
Lemma 2.4 (Primal progress from the smooth adaptable property).
$f(x)-f(x_{+})\geq\gamma\langle\nabla f(x),x-v\rangle-L\gamma^{1+\nu}D_{\phi}(v,x).$ (2.2)
@@
Often the point $v\in P\subset\operatorname{int}\operatorname{dom}\phi$ in Lemma 2.4 is chosen as a Frank–Wolfe vertex, i.e., $v\in\operatorname*{argmax}_{u\in P}\langle\nabla f(x),x-u\rangle$, but other choices, e.g., those arising from away-directions from the away-step FW algorithm, are also possible, as we will see in Section 3.2.
@@
In [garber2015faster] (see also [Garber2020-tr]) the linear convergence of the FW algorithm for convex optimization under the quadratic growth condition, which is a weaker assumption than assuming strong convexity, was established over strongly convex sets and later generalized to uniformly convex sets in [KAP2020] as well as to conditions weaker than quadratic growth in [KDP2018, KDP2018jour]. Local variants of these notions, as necessary, e.g., for the nonconvex case, have been studied in [KAP2021] in the context of FW algorithms. For $f:\mathbb{R}^{n}\to[-\infty,+\infty]$, let $[f\leq\zeta]:=\{x\in\mathbb{R}^{n}\,|\,f(x)\leq\zeta\}$ be a $\zeta$-sublevel set of $f$ for some $\zeta\in\mathbb{R}$.
@@
Definition 2.5 (Hölder Error Bound [Braun2022-ew, Definition 3.25], [alex17sharp, Definition 1.2] and Quadratic Growth Conditions [garber2015faster, Garber2020-tr, Liao2024-vd]).
@@
Let $f:\mathbb{R}^{n}\to(-\infty,+\infty]$ be a proper lower semicontinuous function and $P\subset\mathbb{R}^{n}$ be a compact convex set. Let $\mathcal{X}^{*}\neq\emptyset$ be the set of optimal solutions, i.e., $\mathcal{X}^{*}:=\operatorname*{argmin}_{x\in P}f(x)$, and let $f^{*}=\min_{x\in P}f(x)$ and $\zeta>0$. The function $f$ satisfies the Hölder error bound (HEB) condition on $P$ if there exists constants $q\geq 1$ and $\mu>0$ such that, for any $x\in P$,
$\operatorname{dist}(x,\mathcal{X}^{*})^{q}\leq\frac{q}{\mu}(f(x)-f^{*}).$ (2.3)
Particularly, when $q=2$, it is called the $\mu$-quadratic growth condition. Moreover, the function $f$ is said to be local $\mu$-quadratic growth (with $\zeta$) if there exists a constant $\mu>0$ such that, for any $x\in[f\leq f^{*}+\zeta]\cap P$, (2.3) holds with $q=2$.

### v4

@@ Abstract
We propose Frank–Wolfe (FW) algorithms with an adaptive Bregman step-size strategy for smooth adaptable (also called: relatively smooth) (weakly-) convex functions. This means that the gradient of the objective function is not necessarily Lipschitz continuous, and we only require the smooth adaptable property. Compared with existing FW algorithms, our assumptions are less restrictive. We establish convergence guarantees in various settings, including convergence rates ranging from sublinear to linear, depending on the assumptions for convex and nonconvex objective functions. Assuming that the objective function is weakly convex and satisfies the local quadratic growth condition, we provide both local sublinear and local linear convergence with respect to the primal gap. We also propose a variant of the away-step FW algorithm using Bregman distances over polytopes. We establish faster global convergence (up to a linear rate) for convex optimization under the Hölder error bound condition and local linear convergence for nonconvex optimization under the local quadratic growth condition. Numerical experiments demonstrate that our proposed FW algorithms outperform existing methods.
@@ 1 Introduction
where $f:\mathbb{R}^{n}\to(-\infty,+\infty]$ is a continuously differentiable function and $P\subset\mathbb{R}^{n}$ is a compact convex set. We are interested in both convex and nonconvex $f$ and assume that we have first-order oracle access to $f$, i.e., given $x\in\mathbb{R}^{n}$, we can compute $\nabla f(x)$. The Frank–Wolfe (FW) algorithm is a projection-free first-order method. Instead of requiring access to a projection oracle, the FW algorithm requires only access to a so-called linear minimization oracle (LMO), which for a given linear function $a\in\mathbb{R}^{n}$ computes $y\in\operatorname*{arg\,min}_{v\in P}\langle a,v\rangle$. LMOs are often much cheaper than projection oracles, as shown in Combettes and Pokutta (2021) (see also (Braun et al., 2025, Table 1.1)). Consequently, in practice, FW algorithms are often faster than projected gradient methods even when the projection operation is nontrivial. Additionally, FW algorithms tend to be numerically quite robust and stable due to their affine-invariance and can also be used, e.g., to provide theoretical guarantees for the approximate Carathéodory problem by Combettes and Pokutta (2023).
@@
The FW algorithm was originally proposed by Frank and Wolfe (1956) and was independently rediscovered and extended by Levitin and Polyak (1966) as the conditional gradient method; we will use these terms interchangeably. Canon and Cullum (1968) established an initial lower bound for the rate of the FW algorithm. GuéLat and Marcotte (1986) improved this bound and also provided the first analysis of the away-step FW algorithm by Wolfe (1970). Jaggi (2013) provided a more detailed convergence analysis of the FW algorithm, establishing a new lower bound that demonstrated a trade-off between sparsity and error. Concurrently, Lan (2013) examined the complexity of linear programming-based first-order methods, establishing a similar lower bound. For a comprehensive review of FW algorithms, we refer the interested reader to the survey by Braun et al. (2025) and the brief introduction by Pokutta (2024).
The classical FW algorithm (Algorithm 1) has been improved in two main directions to enhance both performance and convergence. One focuses on refining the step-size rule. The short-step strategy computes $\gamma_{t}$ using the Lipschitz constant $L$ of $\nabla f$. Pedregosa et al. (2020) introduced an adaptive step-size strategy that dynamically estimates $L$. They proved that this strategy is at least as effective as the short-step strategy. It was enhanced by Pokutta (2024) to improve numerical stability.
Another direction is to modify the classical FW algorithm to eliminate the zigzag behavior when approaching the optimal face containing the optimal solution $x^{*}$. This behavior inspired Wolfe (1970) to propose the away-step FW algorithm in 1970 that shortcuts the zigzagging by removing atoms that slow down the iterative sequence. Lacoste-Julien and Jaggi (2015) showed the linear convergence of the away-step FW algorithm. Their analysis [^1] Similar analyses based on the pyramidal width have been developed for other advanced variants, such as the pairwise FW algorithm (Lacoste-Julien and Jaggi, 2015), Wolfe’s algorithm (Lacoste-Julien and Jaggi, 2015), the blended conditional gradient method (Braun et al., 2019), and blended pairwise conditional gradient methods and similar variants (Combettes and Pokutta, 2020; Tsuji et al., 2022). Even Nesterov-style acceleration is possible under weak assumptions (Diakonikolas et al., 2020) building upon this analysis. introduced a geometric constant, the pyramidal width, that measures the conditioning of the polytope $P$, representing the feasible region.
@@
To derive the convergence rate of the above-mentioned FW algorithms, previous studies have typically required $L$-smoothness, i.e., that $\nabla f$ is Lipschitz continuous, and convexity for the objective function $f$, although exceptions exist. [^2] Quite recently, Vyguzov and Stonyakin (2025) proposed a FW algorithm with the Bregman distance, similar to ours, yet it is restricted to convex problems and provides only partial convergence guarantees. These assumptions often narrow the application of FW algorithms. Even if they do not satisfy $L$-smoothness, there are many functions that satisfy the $L$-smooth adaptable property ($L$-smad) with kernel generating distances $\phi$ (see Definition 2.1) by choosing $\phi$ well to match the function $f$. Since the $L$-smad property is consistent with $L$-smoothness when $\phi=\frac{1}{2}\|\cdot\|^{2}$, the class of $L$-smad functions includes the $L$-smooth function class. Other first-order algorithms, such as the Bregman proximal gradient algorithm, are often analyzed under the $L$-smad property (see Appendix A and Bolte et al. (2018); Hanzely et al. (2021); Rebegoldi et al. (2018); Takahashi et al. (2022); Takahashi and Takeda (2025); Yang and Toh (2025)). $L$-smad functions appear in many applications, such as nonnegative linear inverse problems (Bauschke and Borwein, 1997; Takahashi and Takeda, 2025), $\ell_{p}$ loss problems (Kyng et al., 2015; Maddison et al., 2021), phase retrieval (Bolte et al., 2018; Takahashi et al., 2022), nonnegative matrix factorization (NMF) (Mukkamala and Ochs, 2019; Takahashi et al., 2026), and blind deconvolution (Takahashi et al., 2023). Any $\mathcal{C}^{2}$ function is locally $L$-smooth over compact sets, but the resulting Lipschitz constant $L$ can be overly conservative, hindering practical performance. Functions outside $\mathcal{C}^{2}$ may not satisfy $L$-smoothness at all. Moreover, functions in the above-mentioned applications are weakly convex on compact sets. For nonconvex functions, although Lacoste-Julien (2016) established sublinear convergence, little else has been done, and simply tracing existing research does not achieve convergence rates better than sublinear convergence. Then the following question naturally arises:
Is it possible to relax the $L$-smoothness and convexity assumptions often made in FW algorithm studies while still achieving the theoretical guarantee of linear convergence rates?
@@
The answer to the above question is yes. We develop FW algorithms for $L$-smad $f$ and give theoretical guarantees of linear convergence under weaker assumptions (Hölder error bound (HEB) condition with parameter $q\geq 1$ for convex $f$ and the special case with $q=2$ called quadratic growth for nonconvex $f$) than the strong convexity assumption that is often made when linear convergence is shown for FW algorithms. The convexity assumption on $f$ is also weakened to weakly convex, i.e., $f+\frac{\rho}{2}\|\cdot\|^{2}$ is convex with some $\rho>0$ (see, e.g.(Davis et al., 2018, Examples 3.1 and 3.2) and Nurminskii (1975)). Any $\mathcal{C}^{2}$ function on a compact set is weakly convex (Vial, 1983; Wu, 2007) (see Proposition C.1); since $\rho$ is only for the theoretical guarantee and not for the algorithm, there is no need for its estimation, and considering weakly convex $f$ greatly broadens applicability.
@@
Table 1 summarizes our contributions for the various cases that we consider. Relaxing the assumption from $L$-smoothness to the $L$-smad property extends the FW algorithm to a broader range of problems, while making the construction of the FW algorithm and its theoretical guarantee much more difficult. This is because we use the Bregman distance $D_{\phi}$, which is an extension of the Euclidean distance, for the $L$-smad property. While Equation 2.1 shown later is simple with $\nu=1$ when using the Euclidean distance, $\nu>0$ cannot be eliminated for the general Bregman distance; $1+\nu$ represents its scaling exponent. $\nu$ needs to be estimated during the algorithm since its exact value is unknown, and thus, the extension to an $L$-smad $f$ is not straightforward.
Table 1: Our convergence rates are generalizations of existing rates. Moreover, our convergence rate is better than $\mathcal{O}(\epsilon^{-1})$ in some cases. For nonconvex functions, convergence is measured using the Frank–Wolfe gap $\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle\leq\epsilon$, instead of the primal gap $f(x_{t})-f^{*}\leq\epsilon$. Convergence rates for weakly convex optimization hold locally.
Assumptions333ff conv. means the convexity of $f$, IP means $x^{*}\in\operatorname{int}P$, and poly. means that $P$ is a polytope. Convergence rate
FW $f$ conv. $f$ growth IP poly. $L$-smooth $L$-smad ($\nu>0$)
Alg.1 (any step-size) conv. $\mathcal{O}(\epsilon^{-1})$ $\mathcal{O}(\epsilon^{-1/\nu})$ (Thm.D.2)
[…]
away-step FW conv. $q$-HEB $\mathcal{O}(\log\epsilon^{-1})$, $\mathcal{O}(\epsilon^{\frac{2-q}{q}})$44footnotemark: 4 $\mathcal{O}(\log\epsilon^{-1})$, $\mathcal{O}(\epsilon^{\frac{1+\nu-q}{\nu q}})$ (Thm. 4.4)55footnotemark: 5
Alg.1 (any step-size) weak 2-HEB $\mathcal{O}(\epsilon^{-1})$ $\mathcal{O}(\epsilon^{-1/\nu})$ (Thm. E.3)
[…]
away-step FW weak 2-HEB $\mathcal{O}(\log\epsilon^{-1})$ $\mathcal{O}(\log\epsilon^{-1})$, $\mathcal{O}(\epsilon^{\frac{\nu-1}{2\nu}})$ (Thm. 5.4)66footnotemark: 6
Alg.1 (any step-size) $\mathcal{O}(\epsilon^{-2})$ $\mathcal{O}(\epsilon^{-1-1/\nu})$ (Thm. E.1)
@@
  • We establish sublinear convergence for the case of convex $L$-smad functions. Assuming that $f$ is convex and satisfies the HEB condition, and that the minimizer $x^{*}\in\operatorname{int}P$, we provide faster convergence (up to a linear rate). In this setting, our algorithm always converges linearly if the exponent of the HEB condition $q$ equals the scaling exponent of the Bregman distance $1+\nu$, i.e., $q=1+\nu$. [^7] As discussed in Sec. 2, linear convergence with the Euclidean distance is shown in Garber and Hazan (2015) under the quadratic growth over strongly convex sets. This corresponds to the $q=1+\nu$ case with $\nu=1$. For $q>1+\nu$, if $t\leq t_{0}$ with some $t_{0}\in\mathbb{N}$, linear convergence holds; otherwise, $\mathcal{O}(\epsilon^{(1+\nu-q)/\nu q})$, which is faster than existing sublinear rates. A similar argument also applies to the nonconvex case ($q=2$) and the away-step FW algorithm.
@@
  • For nonconvex optimization, assuming that $f$ is weakly convex and satisfies the local quadratic growth condition, we establish local linear convergence. To the best of our knowledge, this is the first FW algorithm with linear convergence guarantees for a certain class of nonconvex optimization problems, albeit for a rather restricted subclass. These results are also new for the standard setting: $L$-smooth $f$. Without weak convexity, we prove global sublinear convergence to a stationary point of Equation 1.1 for nonconvex $L$-smad $f$.
@@ 2 Preliminaries
Let $C$ be a nonempty open convex subset of $\mathbb{R}^{n}$. Let $\operatorname{dist}(x,C):=\inf_{y\in C}\|x-y\|$ and let $\mathcal{C}^{k}$ be the class of $k$-times continuously differentiable functions for $k\geq 0$.
Definition 2.1 (Kernel Generating Distance (Bolte et al., 2018)).
A function $\phi:\mathbb{R}^{n}\to(-\infty,+\infty]$ is called a kernel generating distance associated with $C$ if it satisfies the following conditions: (i) $\phi$ is proper, lower semicontinuous, and convex, with $\operatorname{dom}\phi\subset\operatorname{cl}C$ and $\operatorname{dom}\partial\phi=C$; (ii) $\phi$ is $\mathcal{C}^{1}$ on $\operatorname{int}\operatorname{dom}\phi\equiv C$. We denote the class of kernel generating distances associated with $C$ by $\mathcal{G}(C)$.
Definition 2.2 (Bregman Distance (Bregman, 1967)).
Given $\phi\in\mathcal{G}(C)$, a Bregman distance $D_{\phi}:\operatorname{dom}\phi\times\operatorname{int}\operatorname{dom}\phi\to\mathbb{R}_{+}$ associated with $\phi$ is defined by $D_{\phi}(x,y):=\phi(x)-\phi(y)-\langle\nabla\phi(y),x-y\rangle$.
The Bregman distance $D_{\phi}(x,y)$ measures the proximity between $x\in\operatorname{dom}\phi$ and $y\in\operatorname{int}\operatorname{dom}\phi$. Since $\phi$ is convex, $D_{\phi}(x,y)\geq 0$ holds. Moreover, when $\phi$ is strictly convex, $D_{\phi}(x,y)=0$ holds if and only if $x=y$. However, the Bregman distance is not always symmetric and does not have to satisfy the triangle inequality. See Example B.1. We use $D_{f}$ in Definition 2.2 for nonconvex $f$ instead of convex $\phi$ for brevity. In the remainder of the paper, we assume that, for a given strictly convex $\phi$ and $C$, there exists $\nu>0$ such that the following holds for all $x,y\in\operatorname{int}\operatorname{dom}\phi$ and $\gamma\in[0,1]$ (see Lemma B.3):
@@
We recall the smooth adaptable property, which is a generalization of $L$-smoothness and was introduced by Bauschke et al. (2017); Bolte et al. (2018); Lu et al. (2018). Its characterizations can be found in Appendix B.2, Example B.5, Remark B.6, and (Bauschke et al., 2017, Proposition 1).
Definition 2.3 ($L$-smooth Adaptable Property).
@@
The $L$-smad property is equivalent to the extended descent lemma (Bolte et al., 2018, Lemma 2.1), which implies that the $L$-smad property for $(f,\phi)$ provides upper and lower approximations for $f$ majorized by $\phi$ with $L>0$. For example, $\log x$ is not $L$-smooth on $x>0$ and $\frac{1}{4}x^{4}$ is not $L$-smooth on $\mathbb{R}$. It is the Bregman distance, rather than the Euclidean distance, that allows us to bound the first-order approximation of these functions. See Example B.5 for further examples of $L$-smad pairs. When $(f,\phi)$ is $L$-smad with a strictly convex function $\phi\in\mathcal{G}(C)$, we have Lemma B.7, i.e.,
$f(x)-f(x_{+})\geq\gamma\langle\nabla f(x),x-v\rangle-L\gamma^{1+\nu}D_{\phi}(v,x),$ (2.2)
@@
where $x_{+}=(1-\gamma)x+\gamma v$ with $x,v\in\operatorname{int}\operatorname{dom}\phi$ and $\gamma\in[0,1]$. Often $v\in P\subset\operatorname{int}\operatorname{dom}\phi$ in Lemma B.7 is chosen as a Frank–Wolfe vertex, i.e., $v\in\operatorname*{arg\,max}_{u\in P}\langle\nabla f(x),x-u\rangle$, but other choices, e.g., those arising from away-directions, are also possible, as we will see in Section 3.2.
@@
In Garber and Hazan (2015) (see also Garber (2020)) the linear convergence of the FW algorithm for convex optimization under the quadratic growth condition, which is a weaker assumption than assuming strong convexity, was established over strongly convex sets and later generalized to uniformly convex sets in Kerdreux et al. (2021b) as well as to conditions weaker than quadratic growth in Kerdreux et al. (2019; 2022). Local variants of these notions, as necessary, e.g., for the nonconvex case, have been studied in Kerdreux et al. (2021a) in the context of FW algorithms. For $f:\mathbb{R}^{n}\to[-\infty,+\infty]$, let $f\leq\zeta]:=\{x\in\mathbb{R}^{n}\,|\,f(x)\leq\zeta\}$ be a $\zeta$-sublevel set of $f$ for some $\zeta\in\mathbb{R}$.
@@
Definition 2.4 (Hölder Error Bound (Braun et al., 2025; Roulet and d’Aspremont, 2017) and Quadratic Growth Conditions (Garber and Hazan, 2015; Garber, 2020; Liao et al., 2024)).
@@
Let $f:\mathbb{R}^{n}\to(-\infty,+\infty]$ be a proper lower semicontinuous function and $P\subset\mathbb{R}^{n}$ be a compact convex set. Let $\mathcal{X}^{*}\neq\emptyset$ be the set of optimal solutions, i.e., $\mathcal{X}^{*}:=\operatorname*{arg\,min}_{x\in P}f(x)$, and let $f^{*}=\min_{x\in P}f(x)$ and $\zeta>0$. The function $f$ satisfies the $q$-Hölder error bound (HEB) condition on $P$ if there exist constants $q\geq 1$ and $\mu>0$ such that $\operatorname{dist}(x,\mathcal{X}^{*})^{q}\leq\frac{q}{\mu}(f(x)-f^{*})$ holds for any $x\in P$. In particular, when $q=2$, it is called the $\mu$-quadratic growth condition. Moreover, the function $f$ is said to satisfy local $\mu$-quadratic growth (with $\zeta$) if there exists a constant $\mu>0$ such that $\operatorname{dist}(x,\mathcal{X}^{*})^{2}\leq\frac{2}{\mu}(f(x)-f^{*})$ holds for any $x\in[f\leq f^{*}+\zeta]\cap P$.